### v1

# I’M HOI: Inertia-aware Monocular Capture of 3D Human-Object Interactions

Abstract
We are living in a world surrounded by diverse and “smart” devices with rich modalities of sensing ability. Conveniently capturing the interactions between us humans and these objects remains far-reaching. In this paper, we present I’m-HOI, a monocular scheme to faithfully capture the 3D motions of both the human and object in a novel setting: using a minimal amount of RGB camera and object-mounted Inertial Measurement Unit (IMU). It combines general motion inference and category-aware refinement. For the former, we introduce a holistic human-object tracking method to fuse the IMU signals and the RGB stream and progressively recover the human motions and subsequently the companion object motions. For the latter, we tailor a category-aware motion diffusion model, which is conditioned on both the raw IMU observations and the results from the previous stage under over-parameterization representation. It significantly refines the initial results and generates vivid body, hand, and object motions. Moreover, we contribute a large dataset with ground truth human and object motions, dense RGB inputs, and rich object-mounted IMU measurements. Extensive experiments demonstrate the effectiveness of I’m-HOI under a hybrid capture setting. Our dataset and code will be released to the community at afterjourney00.github.io/IM-HOI.
{strip}
Figure 1: Taking a monocular RGB video and single inertial measurement unit (IMU) sensor recording, our approach, I’m-HOI achieves efficient and robust capture results on challenging and dynamic human-object interaction (HOI) motions like skateboarding.
1 Introduction
Capturing human-object interactions (HOI) is essential to understanding how we humans connect with the surrounding world, with numerous applications in robotics, gaming, or VR/AR. Yet, an accurate and convenient solution remains challenging in the vision community.
For high-fidelity capture of human-object interactions, early high-end solutions [6, 10, 29] require dense cameras, while recent approaches [11, 73, 4, 27] require less RGB or RGBD video inputs (from 3 to 8 views). Yet, such a multi-view setting is still undesirable for consumer-level daily usage. Instead, the monocular method with more handiest captured devices is more attractive. Specifically, most recent methods [82, 85, 104, 107, 86] track the rigid and skeletal motions of objects and humans using a pre-scanned template or parametric model [47] from a single RGB video input. Yet, inherently due to the RGB-setting, they remain vulnerable to depth ambiguity and the occlusion between human and object, especially for handling challenging fast motions like skateboarding. In contrast, the Inertial Measurement Units (IMUs) serve as a rescue to provide motion information that is robust to occlusion. Actually, IMU-based motion capture is widely adopted in both industry [87] and academia [62, 23, 93, 63]. Recent methods [46, 54] further combine monocular RGB video and sparse IMUs, enabling lightweight and robust human motion capture. However, these schemes mostly focus on human-only scenarios and ignore the interacted objects. Moreover, compared to the sometimes tedious requirement of body-worn IMUs, it’s more natural and convenient to attach the IMU sensor to the captured object, since IMUs have been widely integrated into daily objects like phones and smartwatches. Researchers surprisingly pay less attention to capturing human-object interactions with a minimal amount of RGB camera and IMU. The lack of motion data under rich interactions and modalities also constitute barriers to exploring such directions.
In this paper, we propose I’m-HOI – an inertia-aware and monocular approach for robustly tracking both the 3D human and object under challenging interactions (see Fig. 1). In stark contrast to prior arts, I’m-HOI adopts a lightweight and hybrid setting: a minimal amount of RGB camera and object-mounted IMU. Given the expected technological trend of mobile sensing as more and more RGB cameras and IMUs will be integrated into our surrounding devices, we believe that our approach will serve as a viable alternative to traditional human-object motion capture.
In I’m-HOI, our key idea is adopt a two-stage paradigm to make full use of both the object-mounted IMU signals and the RGB stream, which consists of general motion inference and category-aware motion refinement. For the former stage, we introduce holistic human-object tracking in an end-to-end manner. Specifically, we generate human motions via a multi-scale CNN-based network for 3D keypoints, followed by an Inverse Kinematics (IK) optimization layer. To reason the companion object motions, we progressively fuse the human features with IMU measurements via object-orientated mesh alignment feedback. We also adopt a robust optimization to refine the tracked object pose and improve the overlay performance, especially when the object is invisible in the RGB input. For the second refinement stage, we propose to tailor the conditional motion diffusion models [21, 43] for utilizing category-level interaction priors. During training the diffusion model corresponding to a certain object, we treat the tracked motions and the raw IMU measurements from the previous stage as the condition information. We also adopt a novel over-parameterization representation with extra regularization designs to jointly consider the body, object, and especially hand regions during the denoising process. Thus, our refinement stage not only projects the initial human-object motions onto the category-specific motion manifold but also infills possible hand motions for vividly capturing human-object interactions. To train and evaluate our I’m-HOI, we contribute a large multi-modal dataset of human-object interactions, covering 295 interaction sequences with 10 diverse objects in total 892k frames of recording. We also provide ground truth body, hand, and 3D object meshes, with dense RGB inputs and rich object-mounted IMU measurements. To summarize, our main contributions include:
  • We propose a multi-modal method to jointly capture human and object motions from a minimal amount of RGB camera and object-mounted IMU sensor.
  • We adopt an efficient holistic human-object tracking to progressively fuse the motion features, companion with a conditional diffusion model to refine and generate vivid interaction motions.
  • We contribute a large dataset for human-object interactions, with rich RGB/IMU modalities and ground-truth annotations. Our data and model will be disseminated to the community.
Figure 2: The pipeline of I’m-HOI. Assuming video and inertial measurements input, our approach consists of a general interaction motion inference module (Sec. 3.1) and an category-specific interaction diffusion filter (Sec. 3.2) to capture challenging interaction motions.
2 Related Work
Monocular Human-centric Capture.
Human motion capture from single RGB images and videos has achieved tremendous progress [5, 50, 89, 31, 53, 56, 75, 39, 38, 36, 40, 66, 37, 100, 42, 103, 101] since the parametric body model SMPL [47, 68, 57] was released. However, contextual human-object and human-scene interactions (HOIs/HSIs) reconstruction from monocular input is far more challenging. The pioneer work PHOSA [104] proposes a purely optimization-based framework to estimate static human-object spatial arrangements relying on handcrafted contact heuristics, which is unscalable and error-prone to depth ambiguities. Benefited from emerging 3D interaction motions datasets [69, 15, 109, 7, 16, 4, 24, 22, 83, 12, 26], learning-and-optimization work [85, 25] give promising results by modeling human-object relative distance field in data-driven manner and joint post-optimization. The state-of-the-art video-based method VisTracker [86] further incorporates motion infilling techniques [96] to enable space-time coherent tracking. However, these approaches still suffer from unacceptable runtime cost and unsatisfying accuracy under complex interaction scenarios.
Inertial and Multi-modal Motion Capture.
Complementary to vision-based methods, human motion capture using inertial measurement units (IMUs) has also been extensively studied. Previous commercial solutions [87, 52] can capture accurate and detailed motion with dense sensors. Since the exploration of SIP [80], data-driven methods under sparse sensors configuration [23, 93, 94, 28, 84] realize real-time performance, deriving consumer-level products [72]. To mitigate the inherent defects of single-modal systems, multi-modal approaches [18] fuse inertial signals with RGB [61, 62, 79, 48, 78, 81, 13, 20, 110, 49, 30, 46, 54], RGBD [19, 112], ego-view [95], and LiDAR [67] references, achieving balanced local pose estimation and global localization. In this work, we extend multi-sensor fusion strategy to 3D human-object interactions capture, which is validated profitable for both accuracy and efficiency.
Object-specific Interaction Prior.
Human motion priors have been demonstrated crucial for realism of capture and synthesis on multiple modeling methodologies, including predefined kinematic structure [2, 113], GMM [57], GAN [14, 31, 3], VAE [34, 66, 105, 59], MLP [77] and more cutting-edge, diffusion models [21, 71, 76, 70, 97, 32]. Additionally, context-aware human motion synthesis [106, 111, 51] and scene placement generation [91, 92] successfully extract contextual prior knowledge from data. More recently,  [58, 88, 44] model dynamic interaction patterns but ignore object category-level distribution differences. On another track,  [108, 41] focus on interactions with chairs, which is static and lacks diversity. This work aims to learn object category-specific interaction prior to model dynamic interaction distributions between human and diverse objects.
3 Method
We present a new paradigm for 3D human-object interactions capture in a lightweight and hybrid setting: using a minimal amount of RGB camera and object-mounted Inertial Measurement Unit (IMU). As illustrated in Figure 2, we propose a general interaction motion inference module (Sec. 3.1) to jointly recover human-object spatial arrangements in end-to-end manner. An category-specific interaction diffusion filter (Sec. 3.2) is tailored to refine capture results from the former with the learned object category-level prior.
3.1 General Interaction Motion Inference
Existing vision-based methods [85, 86] follow fitting-learning-optimization framework, which, in our observation, is fragile to heavy or prolonged human-object occlusions, inefficient in inference time cost and poor in generalization ability as reported in Sec. 5.2. In contrast, we take object as an extra body joint, and propose to estimate human-object spatial arrangement in a holistic and end-to-end manner. An optional optimization procedure can be implemented to further improve capture accuracy.
Preprocessing.
Given a monocular image sequence $\bm{I}\in\mathbb{R}^{T\times h\times w\times 3}$, we first segment human and object mask $\bm{S}_{h},\bm{S}_{o}\in\mathbb{R}^{T\times h\times w\times 1}$ separately using SAM [35]. Following that, a pre-trained ResNet-34 [17] image encoder is adopted to extract image feature from stacked RGB and object mask. After that, We take the raw inertial rotation $\bm{Q}\in\mathbb{R}^{T\times 6}$, acceleration $\bm{A}\in\mathbb{R}^{T\times 3}$ and normalized object template $\mathcal{O}$, combined with $\bm{I},\bm{S}_{o}$ as our network input. Our approach outputs human shape $\bm{\beta}\in\mathbb{R}^{T\times 10}$ and pose $\bm{\theta}\in\mathbb{R}^{T\times 3N_{J_{b}}}$, object rotation $\bm{R}_{o}\in\mathbb{R}^{T\times 6}$ and translation $\bm{T}_{o}\in\mathbb{R}^{T\times 3}$. Here, $T$ is the sequence length, $h\times w$ is the resolution of images and masks, $N_{J_{b}}=22$ is the number of body joints. We adopt standard SMPL model [47] for human motion representation.
End-to-end Holistic Human-Object Tracking.
We first develop a multi-scale CNN-based network to detect 3D human body joints $\bm{J}\in\mathbb{R}^{T\times 3N_{J_{b}}}$ and object center together. Given the extracted image feature, we feed it into a group of deconvolution layers appended with a final convolution layer to reconstruct 3D keypoint heatmaps. Then we take the expectation of each heatmap to determine 3D keypoint positions [74]. Note that all the keypoints are canonicalized to root-relative representation. Commonly used combination of 3D keypoints and 2D reprojection loss $\mathcal{L}_{\text{kp3d}}+\lambda_{\text{j2d}}\mathcal{L}_{\text{j2d}}$ is taken to train the CNN network, and simultaneously fine-tune the image extractor. Following 3D keypoint estimation, an off-the-shelf pre-trained inverse kinematics layer proposed in [103] is adopted and fine-tuned by $\mathcal{L}_{\text{twist}}$ to recover human motion $\bm{\theta}$ from estimated $\hat{\bm{J}}$. Please refer to [85, 86, 42, 103] for detailed loss functions.
For object tracking, we can get an initially-posed object through $\mathcal{C}(\bm{Q})\mathcal{O}+\hat{\bm{T}_{o}}$ with estimated object translation and raw rotation data from the object-mounted IMU sensor, where $\mathcal{C}(\cdot)$ is the mapping from 6D rotation representation [114] to rotation matrix. In order to eliminate systematic biases in $\bm{Q}$ and correct inaccurate prediction $\hat{\bm{T}_{o}}$ under occlusions, we attach one MLP-based regressor to each of the intermediate multi-scale image feature grid [100, 101], which forms a feedback loop to estimate corrective increment of object motion progressively. In the $i$-th loop, we first uniformly sample $N_{S}=400$ vertices on the posed object mesh $\mathcal{C}(\hat{\bm{R}}_{o}^{(i)})\mathcal{O}+\hat{\bm{T}}_{o}^{(i)}$. Then, we project vertices onto the $i$-th feature map to obtain object mesh-aligned feature, which is subsequently fed into the $i$-th regressor to predict $\Delta\hat{\bm{R}}_{o}^{(i)}$ and $\Delta\hat{\bm{T}}_{o}^{(i)}$. Particularly, $\hat{\bm{R}}_{o}^{(0)}=\mathcal{C}(\bm{Q})$. The training loss of the feedback network group is defined as $\mathcal{L}_{\text{maf}}=\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$, where $\mathcal{L}_{\text{occ-sil}}$ is the occlusion-aware silhouette loss proposed in [104]. We find that better results can be achieved with an augmented silhouette area loss:
$\mathcal{L}_{\text{area}}=\sum_{i=0}^{N_{F}-1}||\sum\mathcal{D}(\hat{\bm{R}}_{o}^{(i)}\mathcal{O}+\hat{\bm{T}}_{o}^{(i)})-\sum\bm{S}_{o}||_{2}^{2},$ (1)
where $\mathcal{D}(\cdot)$ refers to differentiable rendering function [65] and $N_{F}=3$ is the number of feedback iterations. Starting from here, we re-define $\hat{\bm{R}}_{o}=\hat{\bm{R}}_{o}^{N_{F}-1}$ and $\hat{\bm{T}}_{o}=\hat{\bm{T}}_{o}^{N_{F}-1}$.
Overall, our training loss of the end-to-end inference module is:
$\mathcal{L}=\mathcal{L}_{\text{kp3d}}+\lambda_{\text{j2d}}\mathcal{L}_{\text{j2d}}+\mathcal{L}_{\text{twist}}+\mathcal{L}_{\text{maf}}.$ (2)
Robust and Lightweight Optimization.
To improve object tracking accuracy, especially in invisible cases, we propose an optional optimization module. In addition to visual cues, we further constraint object rotation and trajectory to inertial measurements. The energy function is formulated as:
$\mathcal{E}=\mathcal{E}_{\text{visual}}+w_{\text{imu}}\mathcal{E}_{\text{imu}}.$ (3)
Specifically, $\mathcal{E}_{\text{visual}}$ minimizes the discrepancy between the rendering result and the segmentation of object: $\mathcal{E}_{\text{visual}}=\sum||\mathcal{D}(\hat{\bm{R}}_{o}\mathcal{O}+\hat{\bm{T}}_{o})-\bm{S}_{o}||_{2}^{2}$. At the same time, $\mathcal{E}_{\text{imu}}$ regularizes object motion temporally:
$\mathcal{E}_{\text{imu}}=\frac{1}{T-1}\sum_{t=1}^{T-1}$ $||(\hat{\bm{T}}_{o,t-1}+\hat{\bm{T}}_{o,t+1}-2\hat{\bm{T}}_{o,t})-0.5\bm{A}_{t}^{2}||_{2}^{2}$ (4)
$+\frac{1}{T}\sum_{t=0}^{T-1}$ $||\hat{\bm{R}}_{o,t}-\bm{Q}_{t}||_{2}^{2}.$
3.2 Category-specific Interaction Diffusion Filter
For the second refinement stage, a category-specific interaction motion filter is proposed to (i) project capture results from the previous stage onto the manifold; (ii) infill hand motions conditioned on body-object interaction motions.
Interaction Representation.
We propose a novel over-parameterization interaction representation containing human, object motion and raw inertial measurements. At timestamp $t$ and noise level $n$, $\bm{x}_{t}^{n}\in\mathbb{R}^{486}$ consists of body-hand joint positions $\bm{j}_{h,t}\in\mathbb{R}^{156}$ and rotations $\bm{\theta}_{h,t}\in\mathbb{R}^{312}$; object translation $\bm{j}_{o,t}\in\mathbb{R}^{3}$ and rotation $\bm{\theta}_{o,t}\in\mathbb{R}^{6}$; inertial rotation $\bm{q}_{t}\in\mathbb{R}^{6}$ and free acceleration signal $\bm{a}_{t}\in\mathbb{R}^{3}$. We use 6D representation [114] for all the rotation data, and the 52 joints body-hand model SMPL-H [68] is adopted. The target interaction motion is represented as $\bm{x}_{0}$.
[…]
Conditional Diffusion Denoising Process.
Given $\bm{x}_{0}$, forward diffusion process adds Gaussian noise iteratively along an $N$-step Markov chain. For each noising step $n$, the noised interaction motion is drawn from conditional probability distribution determined by a pre-defined schedule $\{\alpha_{n}\}_{n=1}^{N}$:
$q(\bm{x}_{1:T}^{n}|\bm{x}_{1:T}^{n-1})=\mathcal{N}(\sqrt{\alpha_{n}}\bm{x}_{1:T}^{n-1},(1-\alpha_{n})\mathcal{I}).$ (5)
In reverse process, we formulate condition information as a tuple $\bm{c}=(\bm{j}_{h_{b}},\bm{j}_{o},\bm{\theta}_{h_{b}},\bm{\theta}_{o},\bm{q},\bm{a})\in\mathbb{R}^{216}$ and concatenate it with masked hand motion $\bm{m}=(\bm{j}_{h_{h}},\bm{\theta}_{h_{h}})\in\mathbb{R}^{270}$, where $\bm{j}_{h_{b}}\in\mathbb{R}^{66}$, $\bm{\theta}_{h_{b}}\in\mathbb{R}^{132}$ means body-only joint positions and rotations. We follow [64] to predict $\bm{x}_{0}$ itself as $\hat{\bm{x}}_{\phi}(\bm{x}_{n},n,\bm{c})$, where $\phi$ represents the parameters of neural network. The training loss is $L1$-norm simple objective [21, 43]:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{\bm{x}_{0},n}||\hat{\bm{x}}_{\phi}(\bm{x}_{n},n,\bm{c})-\bm{x}_{0}||_{1}.$ (6)
Inspired by [66], components inside of the over-parameterization representation have mutual constraints. We accordingly introduce four regularization terms:
$\mathcal{L}_{\text{reg}}=\lambda_{\text{off}}\mathcal{L}_{\text{off}}+\lambda_{\text{vel}}\mathcal{L}_{\text{vel}}+\lambda_{\text{consist}}\mathcal{L}_{\text{consist}}+\lambda_{\text{imu}}\mathcal{L}_{\text{imu}}.$ (7)
Specifically, $\mathcal{L}_{\text{off}}$ enforces the predicted object center that lies in a small region determined by the offsets relative to 52 body-hand joints as:
$\mathcal{L}_{\text{off}}=\sum_{i=0}^{N_{J}}||(\hat{\bm{j}}_{o}-\hat{\bm{j}}_{h}^{(i)})-(\bm{j}_{o}-\bm{j}_{h}^{(i)})||_{1}.$ (8)
We then constraint the reproduced human body-hand joints consistent with the body model skinned by predicted joint rotations:
$\mathcal{L}_{\text{consist}}=\frac{1}{T}\sum_{t=0}^{T-1}||\hat{\bm{j}}_{h,t}-\mathcal{J}(\mathcal{M}(\hat{\bm{\beta}},\hat{\bm{\theta}}_{h,t}))||_{1},$ (9)
where $\mathcal{M}(\cdot)$ refers to forward function of SMPL-H and $\mathcal{J}(\cdot)$ is the joint regressor. In order to temporally smooth human motion, the velocity term $\mathcal{L}_{\text{vel}}$ is formulated as:
$\mathcal{L}_{\text{vel}}=\frac{1}{T-1}\sum_{t=1}^{T-1}||(\hat{\bm{j}}_{t}-\hat{\bm{j}}_{t-1})-(\bm{j}_{t}-\bm{j}_{t-1})||_{1},$ (10)
where $\bm{j}_{t}=[\bm{j}_{h,t},\bm{j}_{o,t}]$. Finally, $\mathcal{L}_{\text{imu}}$ guides the generated object poses and the trajectory conform with IMU measurements, which improves robustness under invisible scenarios:
$\mathcal{L}_{\text{imu}}=\mathcal{L}_{\text{rot}}+\mathcal{L}_{\text{acc}}.$ (11)
Wherein, object rotation is directly regularized by:
$\mathcal{L}_{\text{rot}}=||\hat{\bm{\theta}}_{o}-\bm{q}||_{1},$ (12)
and the trajectory is constrained to acceleration through:
$\mathcal{L}_{\text{acc}}=||(\hat{\bm{j}}_{o,t}-\hat{\bm{j}}_{o,t-1}+\frac{\bm{a}_{t}\tau^{2}}{2})-(\bm{j}_{o,t+1}-\bm{j}_{o,t})||_{1},$ (13)
where $\tau$ is the time interval between two consecutive frames. It is noted that, related work [86] simulates such second-order constraints in a pseudo manner [99] to eliminate mutation of first-order signals. In contrast, we incorporate acceleration explicitly [46, 67].
[…]
CD (per-frame) CD ($10s$) CD (per-frame) CD ($10s$) CD (per-frame) CD ($10s$)
Method smpl object smpl object smpl object smpl object smpl object smpl object
PHOSA [104] 29.20 20.26 41.26 56.80 12.86 26.90 27.01 59.08 11.20 20.57 24.16 43.06
CHORE [85] 14.20 16.81 24.32 31.76 5.55 10.02 18.33 20.32 7.12 12.59 16.11 21.05
VisTracker [86] 19.96 23.28 17.02 18.10 5.25 8.04 7.81 8.49 6.76 10.32 9.35 11.38
Ours 6.50 6.93 5.36 8.53 5.26 7.43 5.65 4.82 5.66 8.92 5.81 7.14
Table 1: Quantitative comparison with several baselines on both human and object tracking accuracy.
4 Dataset
[…]
Capture Preparations.
[…]
Data Processing.
Given multi-view videos, we reproduced human motions in SMPL-H format [68] using an open-source toolbox [1]. In order to track accurate object pose in 3D scene, we first manually annotated single key-frame segmentation in all views, and broadcasted to the whole sequence [90, 35, 8, 45]. Subsequently, we optimized euclidean transformations, which can exactly define object motions, by fitting reprojected silhouette to multi-view masks. For object geometries, we utilized a public application [60] to get 3D scan templates. In terms of inertial signals, we adopted primitive rotation data $\bm{R}_{s}$ in matrix form and transformed raw acceleration $\bm{a}_{raw}$ in sensor coordinate to free acceleration $\bm{a}_{free}$ in global coordinate through $\bm{a}_{free}=\bm{R}_{s}\bm{a}_{raw}-\bm{g}$, where $\bm{g}=[0,0,9.81]^{\mathrm{T}}$ is the gravitational acceleration.
Figure 5: Qualitative comparison results. I’m-HOI performs better than baselines and generalizes well to new datasets.
5 Experiments
In this section, we first introduce the datasets and metrics used for training and evaluation. Then we comprehensively compare our approach to baseline methods. We also extensively conduct ablation studies to demonstrate the effectiveness of pivot components in our network design and the necessity of IMU modality.
5.1 Datasets and Evaluation Metrics
[…]
Evaluation Metrics.
  • Per-frame Chamfer Distance ($cm$) [85] computes the chamfer distance between predicted human and object mesh with the ground truth respectively after holistic Procrustes alignment for every single frame.
  • Sliding Window Chamfer Distance ($cm$) [86] computes the chamfer distance in the same way but performing holistic Procrustes alignment on combined mesh of 10-second results with the ground truth.
5.2 Comparison
Results.
[…]
Generalization.
In order to validate the generalization ability, we apply purely optimization-based method PHOSA [104], learning-and-optimization methods CHORE[85], VisTracker[86] and our approach trained on BEHAVE[4] and InterCap[24] dataset to HODome[102] and CHAIRS[26] dataset. As shown in Table 2, I’m-HOI generalizes better than baselines by a large margin and achieves more balanced performance between per-frame and sequential results. The illustration of Figure 5 shows that I’m-HOI can adapt to various scene backgrounds and objects.
Runtime Cost.
[…]
CD (per-frame) CD ($10s$)
Dataset Method smpl object smpl object
HODome [102] PHOSA [104] 34.41 29.70 60.15 54.98
CHORE [85] 23.18 16.18 43.35 31.64
VisTracker [86] 11.87 19.86 32.77 34.53
Ours 8.19 9.05 12.07 15.31
CHAIRS [26] PHOSA [104] 35.26 28.35 43.17 37.67
CHORE [85] 19.10 36.13 16.71 52.95
VisTracker [86] 17.42 23.31 15.23 16.85
Ours 9.55 9.91 6.34 7.85
Table 2: Quantitative evaluations of generalization ability. I’m-HOI outperforms baselines by a large margin on new data.
5.3 Ablation Study
[…]
Network Architecture.
Table 3 shows the performance of models with or without the mesh alignment feedback (maf.), optimization module (optim.) and diffusion filter (filter.). It is demonstrated that maf. improves per-frame object tracking results and optim. brings better temporal consistency. Moreover, filter. further corrects human-object spatial arrangements onto the learned interaction manifold. Compared with naive implementation, the full pipeline of I’m-HOI performs 4 times better. Figure 6 illustrates that inaccurate prediction is corrected progressively when maf. and optim. are applied. Also, the hand motion generated from filter. makes the capture result more vivid and realistic.
Module CD (per-frame) CD ($10s$)
maf. optim. filter. smpl object smpl object
8.42 15.12 13.88 27.85
8.02 9.73 10.31 19.87
7.16 7.42 7.25 10.75
6.50 6.93 5.36 8.53
Table 3: Quantitative evaluations of network architecture design.
Figure 6: Qualitative evaluation of our network architecture. The figure illustrates the effectiveness of each key design.
Input Modality.
We experiment on several different baselines with IMU modality input by adding extra inertial optimization term described in Equation 4 to their pipelines. Qualitative results illustrated in Figure 7 clearly exhibit the improvements on object pose estimation of baselines after introducing IMU modality, compared with Figure 5. Quantitative performance reported in Table 4 shows decent increase of baselines’ performance, compared with the statistics in Table 1. Additionally, we observe that our approach achieves better results when the IMU modality is involved, especially for object tracking. Moreover, Table 4 indicates that naively incorporating IMU modality input within baselines is unable to maximize its advantages, which further verifies the effectiveness of our network design.
5.4 Limitation
As the first trial to explore challenging 3D human-object interactions capture using a minimal amount of RGB camera and object-mounted IMU sensor, the proposed I’m-HOI still owns limitations as follows. First, our method relies on pre-scanned object template and manual manipulation of sensor-template coordinate alignment. Our method is also restricted to rigid object tracking. It’s a promising direction to extend the interesting task to articulated or even deformable objects in template-free framework.
CD (per-frame) CD ($10s$)
Modality smpl object smpl object
PHOSA+imu 28.41 18.60 39.19 38.44
CHORE+imu 12.98 11.92 22.09 23.31
VisTracker+imu 15.87 14.61 14.68 12.82
Ours w/o imu 10.58 14.49 6.55 15.23
Ours 6.50 6.93 5.36 8.53
Table 4: Quantitative evaluations on input modality configurations.
Figure 7: Qualitative evaluation of the IMU modality. This figure shows the importance of inertial measurements input.
6 Conclusion
We have presented a novel and monocular scheme to faithfully capture the 3D motions of human-object interactions, using a minimal amount of RGB camera and object-mounted IMU. Our general motion inference stage progressively fuses the IMU signals and the RGB stream via holistic and end-to-end tracking, which efficiently recovers the human motions and subsequently the companion object motions via mesh alignment feedback. Our category-aware motion diffusion further treats the previous results as conditions and jointly considers the body, object, and especially hand regions during the denoising process with an over-parameterization representation. It encodes category-aware motion priors, so as to significantly improve the tracking accuracy and generate vivid hand motions. Our experimental results demonstrate the effectiveness of I’m-HOI for faithfully capturing human and object motions in a lightweight setting. As more and more sensors like RGB cameras or IMUs will be integrated into our surrounding world, we believe that our approach and dataset will serve as a critical step towards hybrid human-object motion capture, with many potential applications in robotics, embodied AI, or VR/AR.
Supplementary Material
[…]
Appendix A Implementation Details
A.1 General Interaction Motion Inference
Network Architecture.
We use the pre-trained ResNet-34 [17] $f^{\text{enc}}:\mathbb{R}^{H\times W\times 4}\mapsto\mathbb{R}^{8\times 8\times 512}$ to extract image feature, where $H=W=256$. Then we use 3 stacked deconvolution layers to construct feature pyramid, each of which $f^{\text{deconv}}_{i}:\mathbb{R}^{H_{\text{in}}\times W_{\text{in}}\times C_{\text{in}}}\mapsto\mathbb{R}^{2H_{\text{in}}\times 2W_{\text{in}}\times C_{\text{out}}}$ takes input feature with resolution $H_{\text{in}}=W_{\text{in}}=8,16,32$ and channel $C_{\text{in}}=512,256,256$ respectively, and outputs 256-channel 2 times upsampled feature map. Each deconvolution layer is followed by a Batch Normalization layer and ReLu activation function. For every single intermediate feature map, one regressor $f^{\text{reg}}_{i}:\mathbb{R}^{H_{\text{in}}\times W_{\text{in}}\times 256}\mapsto\mathbb{R}^{6+3}$ is tailored to embed mesh-aligned feature to $\mathbb{R}^{2000}$ and concatenates it with $\hat{\bm{R}}_{o}^{(i-1)},\hat{\bm{T}}_{o}^{(i-1)}$ to predict $\Delta\hat{\bm{R}}_{o}^{(i)},\Delta\hat{\bm{T}}_{o}^{(i)}$. All regressors consist of two hidden Linear layers with dimension $1024$ and two output Linear layers predicting delta rotation and translation separately. Dropout with $0.5$ probability are inserted between each two adjacent Linear layers.
Training.
The proposed $f^{\text{enc}},\{f^{\text{deconv}}_{i}\}_{i=1}^{3},\{f^{\text{reg}}_{i}\}_{i=1}^{3}$ are trained end-to-end with the inverse kinematics layer, supervised by $\mathcal{L}=\mathcal{L}_{\text{kp3d}}+\lambda_{\text{j2d}}\mathcal{L}_{\text{j2d}}+\mathcal{L}_{\text{twist}}+\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$. Particularly, the object-oriented mesh alignment feedback loss $\mathcal{L}_{\text{maf}}=\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$ is added after 55 training epochs. The loss weights are: $\lambda_{\text{j2d}}=1\times 10^{-9},\lambda_{\text{occ-sil}}=1\times 10^{-6},\lambda_{\text{area}}=2\times 10^{-7}$. The model is trained for 190 epochs on 6 NVIDIA GeForce RTX 3090 GPUs. In each epoch, we randomly downsample images 8$\times$ to train and the training batch size is set to 8.
Optimization.
[…]
A.2 Category-specific Motion Diffusion Filter
Network Architecture.
We use 4 transformer encoder-only layers with 4 attention heads to learn category-specific human-object interaction manifold. The model dimension $D_{\text{model}}=1024$ and key, value dimension $D_{\text{key}}=D_{\text{value}}=512$. We take standard $N=1000$ denoising steps and sinusoidal positional encoding function. Different from [43] in which the condition is exactly a part of the target motion, we take results from previous stage together with raw IMU measurements as condition to model the transformation from prediction distribution to the true manifold.
Training.
We first warm up the diffusion model on our whole training dataset only with simple objective function for 100 epochs. After that, we train the model on category-specific data with designed regularization terms $\mathcal{L}_{\text{consist}},\mathcal{L}_{\text{vel}}$ and $\mathcal{L}_{\text{imu}}$ to implicitly model particular interaction patterns. The regularization term weights are $\lambda_{\text{off}}=\lambda_{\text{vel}}=\lambda_{\text{consist}}=1,\lambda_{\text{imu}}=100$. More detailed, we apply $\mathcal{L}_{\text{off}}$ and $\mathcal{L}_{\text{vel}}$ for 35 epochs before adding $\mathcal{L}_{\text{consist}}$ and $\mathcal{L}_{\text{imu}}$. For better generation, we maintain an exponential moving average (EMA) version model during training, which is updated per 10 epochs with decay rate $0.995$. We also enable automatic mixed precision (AMP) to accelerate training process. The model is trained for 55 epochs on single NVIDIA GeForce RTX 3090 GPU. The training batch size is set to 128.
[Uncaptioned image] Holdhandle Hit Lefthand Carry [Uncaptioned image] Ollie
Holdhead Hit [Uncaptioned image] Lefthand Push Kickflip
Lefthand Swing Lift Grind
Midpart Rotate Putdown Pickup Manual
Pickup Putdown Ride Play Heelflip
Righthand Swing Righthand Carry Pop Shove-it
Rub Righthand Push Nollie
Throw Catch Twohands Carry Varial Kickflip
Twoends Rotate suitcase Twohands Push McTwist
baseball bat Twohands Swing Twohands Pull skateboard Darkslide
[Uncaptioned image] Left Biceps [Uncaptioned image] Forward Swing [Uncaptioned image] Forehand [Uncaptioned image]
Left Lunges Backward Swing Backhand Hold
Left Triceps Snatch Volley Stir
Right Biceps Turkish Get-up Overhead Smash Shake
Right Lunges Goblet Squat Slice Flip
dumbbell Right Triceps kettlebell Windmill tennis racket Drop Shot pan
[Uncaptioned image] Drive Sit [Uncaptioned image] Sweep
Putt [Uncaptioned image] Lean Push
Chip Adjust Pull
Pitch Swivel Twist
Sand Shot Recline Store
Fade Rest Tap
Hook Clean Tilt
Draw Lift Lift
Grip chair Rock Grip
golf club Slice Kick broom Maintain
[…]
Appendix B Data Preparation Details
[…]
Temporal Synchronization.
To synchronize RGB data and IMU measurements, we asked the performer to wear an extra IMU sensor on his/her ankle part and make a takeoff move at the beginning of each interaction split. We automatically identified the moment the performer falling to the ground through gravitational acceleration mutation from IMU signals, and manually annotated that in RGB sequences.
Spatial Alignment.
To mitigate spatial misalignment between camera and IMU, we conducted spatial alignment once per ten minutes. Specifically, in our multi-modal multi-sensor system, there exists multiple coordinate frames, including $\{\mathcal{F}_{C_{i}}\}_{i=0}^{31}$ for cameras, $\mathcal{F}_{W}$ for world and $\mathcal{F}_{I}$ for inertia. Since the transformation $\mathcal{T}_{W\rightarrow C_{i}}\in\mathcal{SO}(3)$ from $\mathcal{F}_{W}$ to $\mathcal{F}_{C_{i}}$ is easy to obtain through off-the-shelf toolbox, our goal is to calibrate the transformation $\mathcal{T}_{I\rightarrow W}\in\mathcal{SO}(3)$ from $\mathcal{F}_{I}$ to $\mathcal{F}_{W}$.
In implementation, we capture the global orientation $\{\bm{R}_{t}^{W}\in\mathcal{SO}(3)\}_{t=0}^{T-1}$ of a performer who wears an IMU sensor on his/her waist part and circles around in $\mathcal{F}_{W}$ using [1], and record the inertial rotation measurement $\{\bm{R}_{t}^{I}\in\mathcal{SO}(3)\}_{t=0}^{T-1}$ in $\mathcal{F}_{I}$. Suppose that the IMU sensor is relatively fixed to the performer, we can construct the following equation:
$\mathcal{T}_{I\rightarrow W}\bm{R}_{t}^{I}(\mathcal{T}_{I\rightarrow W}\bm{R}_{t+s}^{I})^{-1}=\bm{R}_{t}^{W}(\bm{R}_{t+s}^{W})^{-1},$ (14)
where $s=5$ is the stride. Let $\bm{B}_{t}=\bm{R}_{t}^{I}(\bm{R}_{t+s}^{I})^{-1}$ and $\bm{A}_{t}=-\bm{R}_{t}^{W}(\bm{R}_{t+s}^{W})^{-1}$, we can reformulate Equation 14 as:
$\bm{A}_{t}\mathcal{T}_{I\rightarrow W}+\mathcal{T}_{I\rightarrow W}\bm{B}_{t}=\bm{0},$ (15)
which is a Sylvester equation. To solve this equation, both analytical [55] and iterative optimization methods [33] can be used.
B.2 Acceleration Data Processing
Normalization on Real Data.
Assume that all the objects are rigid and have uniform rotational inertia, the centroid of each object is equivalent to its geometry center. However, it’s impractical to mount IMU sensor exactly on the center that is probably inside of the object. Hence, extra linear acceleration can be generated even from purely rotational motion, which is undesirable. To eliminate such noises, we first fix a mounting point for each object and manually measure the directional offset $\vec{\bm{r}}$ from center to that point in mesh processing software [9].
Figure 8: Illustration of why extra linear acceleration occurs.
According to recorded angular velocity $\vec{\bm{w}}$, the extra linear velocity generated from rotation is $\vec{\bm{v}}=\vec{\bm{w}}\times\vec{\bm{r}}$, and we can determine the extra linear acceleration: $\delta\bm{a}_{t}=\frac{\bm{v}_{t}-\bm{v}_{t-\Delta t}}{\Delta t}$. Finally, the normalized acceleration data can be achieved by subtracting $\delta\bm{a}_{t}$ from the raw measurements.
Simulation on Synthetic Data.
We also simulate synthetic IMU data based on ground-truth object motion annotations of [4, 24, 102, 26]. In particular, for inertial acceleration data, we follow [93, 46, 67] to calculate second-order difference of object translation:
$\bm{a}_{t}=\frac{\bm{T}_{o,t-n}+\bm{T}_{o,t+n}-2\bm{T}_{o,t}}{(n\tau)^{2}},$ (16)
where $n=4$ is the smoothing factor for better approximation to real acceleration, and $\tau=\frac{1}{\text{fps}}$ is the time interval between two consecutive frames.
B.3 Dataset Statistics
[…]
Figure 9: Attributes comparison between different datasets.
Appendix C More Experiments
C.1 More Results
In former sections, we show the robustness of I’m-HOI under severe occlusions. Here, we further exhibit sequential capture results of I’m-HOI to demonstrate spatial-temporal coherence. Figure 10 illustrates that our approach captures accurate and consistent human-object spatial arrangements in a time window, which validates that our proposed network learns reasonable interaction distributions and recognizes continuous interaction behaviors from input with hybrid modality.
C.2 More Comparisons
We also present more qualitative comparisons on sequential capture results with baselines in Figure 11 and Figure 12. It can be observed that even within an extremely short time interval (around 0.07 seconds), jittery object tracking results occur in image-based baselines [104, 85] which model the static interactions but ignore temporal information. While, video-based method [86] gives temporally consistent but erroneous predictions without inertial measurements, especially for the tracking of object rotational motion. In stark contrast, I’m-HOI makes use of both visual cues and IMU signals, cooperating with design of object-oriented mesh alignment feedback and category-specific interaction prior, which benefits to consistent and correct results.
C.3 Ablation on Regularization Terms
CD (per-frame) CD ($10s$)
Regularization terms smpl object smpl object
w/o $\mathcal{L}_{\text{off}}$ 7.07 7.31 7.59 9.17
w/o $\mathcal{L}_{\text{vel}}$ 7.06 7.62 7.62 10.66
w/o $\mathcal{L}_{\text{consist}}$ 6.29 6.98 6.30 9.01
w/o $\mathcal{L}_{\text{imu}}$ 7.10 7.61 7.87 10.94
Ours 6.50 6.93 5.36 8.53
Table 6: Quantitative evaluations on regularization terms.
To further evaluate the effectiveness of regularization terms in interaction diffusion filter training, we compare the performance of full model with downgraded versions that discard each term. As reported in Table 6, $\mathcal{L}_{\text{off}}$, restricting object to a more specific and precise region, improves accuracy. $\mathcal{L}_{\text{consist}}$ regularizes predicted joint rotations to comply with detected 3D joints after forward kinematics, which prevents overfitting to pseudo ground-truth annotations. Both $\mathcal{L}_{\text{vel}}$ and $\mathcal{L}_{\text{imu}}$ improves performance in temporal dimension. However, only applying $\mathcal{L}_{\text{vel}}$ can lead to oversmooth results since the physical dynamics is lost. Incorporating second-order supervision $\mathcal{L}_{\text{imu}}$ is verified profitable for not only smooth but also physically plausible capturing.
[…]
Figure 11: More qualitative comparisons. I’m-HOI performs better than baselines on sequential data.
Figure 12: More qualitative comparisons. I’m-HOI performs better than baselines on sequential data.